# Twisted Superfields

### v7

@@ 5.8 Scalars
As discussed in section 3, in an $N=2$ Super Yang-Mills (SYM) theory where the adjoint superfield has a tree-level mass, the adjoint scalars become massless color monopoles. The monopoles condense and cause confinement. Since the $m=1$ sector of this theory has a very similar moduli space, it was proposed in section 3 that the octet of $m=1$ SU(3) adjoint scalars of this theory become monopoles, condense, and cause color confinement in the same way.
The six $m=1$ scalar triplets are leptoquarks since they connect gaugino quarks with leptons (and possibly also adjoint-representation quarks with leptons). The model puts no constraint on the scalar masses $m_{1F}$ and $\tilde{m}_{1F}$. It is therefore assumed that these masses are larger than the current lower limits from leptoquark searches. Having large masses for these scalars also strengthens the above confinement argument, since an effective theory below the leptoquark masses has an $m=1$ moduli space that is even more similar to that of $N=2$ SYM. If the flavor-1 leptoquarks have masses of around 10 TeV, they could be at least partly responsible for the non-resonant di-electron anomaly discussed in [5].
The model requires the masses $\tilde{m}_{11}$ and $m_{22}$ of the scalar triplets $\tilde{\phi}_{21}$ and $\phi_{22}$ to be larger than the electroweak scale. The neutral scalar $(\phi_{22})_{2}$ has the right structure from its gauge interactions to generate di-photon, di-Z, di-W, and $b\bar{b}$ events. If $(\phi_{22})_{2}$ had a mass of $\sim$650 GeV it could be responsible for the excesses seen in these events. Quantum effects could generate larger masses of $\sim$850-950 GeV for the charged and neutral scalars $(\phi_{22})_{1}$ and $(\phi_{22})_{3}$. These scalars could decay into two jets and be responsible for the di-jet excesses seen at these masses by CMS and ATLAS.
It was suggested above in the fourth charged lepton section that the charged $(\tilde{\phi}_{21})_{2}$ scalar may have a mass of 2.1 or 2.9 TeV. It is possible that the neutral and charged scalars $(\tilde{\phi}_{21})_{1}$ and $(\tilde{\phi}_{21})_{3}$ may have masses of $\sim$3.5 TeV. From the quantum-generated superpotential coupling of $(\tilde{\phi}_{21})_{1}(f_{22})_{2}(\phi_{22})_{3}$ and the vev $\bar{\tilde{\phi}}_{21}$, the $(\tilde{\phi}_{21})_{1}$ scalar could decay to $(\phi_{22})_{3}(\phi_{22})_{3}^{\dagger}$. It could then be responsible for the di-di-jet excess seen at the LHC. Quantum couplings could also cause the charged scalar $(\tilde{\phi}_{21})_{3}$ to decay to $tb$, generating the excess seen in those data. The anomalies described in this paragraph and the one before it are discussed (and cited) in [5].
The observed Higgs Boson accounts for the SU(2) doublet within the triplet $\phi_{21}$. It was proposed in the fourth lepton section above that the same mechanism that gives a large mass to the Omega lepton also gives a similar mass to the remaining scalar in the triplet: $(\phi_{21})_{3}$.
The $W$-scale mass from eq (2.8) together with radiative corrections could generate a mass of around 150 GeV for scalars in the $\tilde{\phi}_{22}$ triplet. Similarly, quantum effects could generate masses of around 95 GeV for the SU(2) adjoint triplet scalars within the $\varphi_{2}$ octet. If so (or if the masses were reversed), these scalars could potentially be responsible for the remaining Higgs-like resonant signals discussed in [5].
A superpotential coupling $\tilde{\Gamma}_{211}$ together with the Higgs vev $\bar{\phi}_{21}$, generates a mass for the charged scalar $(\varphi_{2})_{31}$. This mass is assumed to be larger than that of the $W$ boson. Similarly, a large mass is assumed to be generated for the charged scalar: $(\varphi_{2})_{13}$.
The scalars that have not yet been discussed in this section are $(\varphi_{2})_{32}$, $(\varphi_{2})_{23}$ the scalar in $\varphi_{2}$ proportional to $t^{8}$ and the two adjoint singlet scalars. It is assumed that the first of these has a small mass. It was discussed above how this small mass could help this model to reproduce experimental $Z$ boson width measurements. The last three scalars mentioned above do not interact with any gauge bosons below the unification scale. One combination of these scalars ($(\varphi_{2})_{33}$) also has no (or very little) interaction with any sub-unification-scale fermions or scalars. For that reason, it is an interesting dark matter candidate.
@@ Discussion
In sections 1, 2, 4 and the appendix, a model is presented that has a number of attractive theoretical features. For example, it is holomorphic, invariant to local superspace gauge transformations, supports coupling constant unification, and is similar at the unification scale to a theory that has been shown to be free of quadratic divergences to at least two loops.
@@
But even without that work, the model provides a couple of interesting explanations and makes a number of predictions. The model provides explanations for the mechanisms of confinement and neutrino oscillations. The model predicts a fourth charged lepton, a seventh quark (without an eighth), a $Z^{\prime}$ boson, right-chiral neutrinos with masses of a few GeV, right-chiral quark interactions with the W boson, and dozens of additional scalar particles. So far, these predictions do not appear to be ruled out by existing data; in fact, they could provide new physics explanations for many of the anomalies discussed in [5].
@@ Appendix A Quantum Effective potential
The U(3)xU(3) theory of this paper has Abelian groups, so it does not appear to be asymptotically free. However, in this appendix it is implicitly assumed that this theory is an effective theory of a more general asymptotically free theory, applicable above the unification scale. With that assumption in hand, this appendix makes arguments similar to those used for SQCD and proposes general features of nonperturbative superpotential terms.
@@
Since they are in the (3,3*) representations, the gauginos and the fermions from the twisted adjoint superfield contribute zero to both $T^{0}_{V_{R}+}$ and $T^{0}_{A_{R}+}$. In addition, for the second term of (A.4) they generate a factor of $f^{bad}d^{cad}$ which vanishes. Therefore, those fermions do not contribute to the axial anomaly.
@@
The anomaly-generating axial U(1) charges of the superfields $Q_{mF}$, $\tilde{Q}_{mF}$, and $\Phi$ are 1, 1, and 0, respectively. Therefore, the following factors are invariant to the U(3)xU(3) local gauge symmetry and obey the axial selection rule:
@@
where $n$ is a non-negative integer. In the second “meson” configuration above, the flavors are restricted by the local U(1) “+” gauge symmetry to either $F=F^{\prime}=1$ or else $F,F^{\prime}\in\left\{2,3\right\}$.
@@
The following determinants are invariant to the 2/3 global flavor symmetry:
$\Delta_{m_{1}m_{2}m_{3}m_{4}}^{\left(pq\right)}=M_{m_{1}m_{2}22}^{\left(p\right)}M_{m_{3}m_{4}33}^{\left(q\right)}-M_{m_{1}m_{2}23}^{\left(p\right)}M_{m_{3}m_{4}32}^{\left(q\right)}$ (A.7)
Following symmetry arguments similar to those used for SQCD, it is speculated that quantum interactions will generate low energy effective superpotential terms with mass dimension 3 and positive powers of $\Lambda$ that are constructed from the above determinants as well as factors of $M^{\left(n\right)}_{mm^{\prime}11}$, ${\rm Tr}\left(\Phi^{r}\right)$, and other gauge invariant factors such as those mentioned later in this appendix.
@@
$O_{0}=\left({\rm Tr}\left(\Phi^{2}\right)\right)^{2}\left(\sum_{m_{i}}M^{\left(0\right)}_{m_{1}m_{2}11}\Delta_{m_{3}m_{4}m_{5}m_{6}}^{\left(00\right)}\right)^{-1/3}$ (A.8)
where the sum over $m_{i}$ is over all permutations of the $m_{i}$ values being $1$ or $2$. The scalar potential can be derived from the superpotential terms like the one above by taking derivatives to extract “f terms” proportional to $\theta^{2}f$ and inserting them into the f-term part of eq (2.1). The quantum vacuum can be found by minimizing the vev of the scalar potential after including these quantum contributions.
In this process, vevs of scalar potential terms derived from $O_{0}$ will involve terms proportional to
@@
In section 2, it was noted that if the model includes large tree-level masses $\tilde{m}_{21}$ and $m_{22}$, then the vevs $\bar{\tilde{\phi}}_{21}$ and $\bar{\phi}_{22}$ vanish clasically. However, in order to stabilize expressions such as the one above, it is assumed that $\bar{\tilde{\phi}}_{21}$ and $\bar{\phi}_{22}$ acquire small vevs quantum mechanically. This is mentioned in section 3. As shown below, those quantum-generated small vevs lead to the very large right-chiral neutrino Majorana mass that allows the model to reproduce observed neutrino masses and mixing.
The classical scalar potential only restricts the difference $\bar{\phi}_{23}^{2}-\bar{\tilde{\phi}}_{23}^{2}$. However, it can be seen from eq (A.9) that the quantum scalar potential will try to make each of these vevs go to infinity while maintaining the difference. A mechanism like that is what causes Supersymmetric QCD with fewer flavors than colors to not have a vacuum solution. But in this theory, a nonzero superpotential coupling $\Gamma_{232}$ can change the result. At tree level, this coupling generates a contribution to the scalar potential of $\Gamma_{232}\bar{\tilde{\phi}}_{23}\bar{\varphi}_{2}\bar{\phi}_{22}$. This contribution vanishes classically, since $\bar{\phi}_{22}$ vanishes classically. But upon quantum generation of $\bar{\phi}_{22}$, this contribution no longer vanishes. Instead, it provides a small counterbalancing effect that stops $\bar{\tilde{\phi}}_{23}^{2}$ from becoming infinite.
It is assumed that minimizing the quantum scalar potential generates the adjoint vevs in $\left\langle\varphi_{2}\right\rangle$ of eq (3.4). As an example for how this works, if $\Phi$ only appeared in powers of ${\rm Tr}\left(\Phi^{2}\right)$ in the numerator, then minimization of the scalar potential for the vacuum would try to make that trace vanish by imposing ${\rm tr}\left\langle\varphi_{2}^{2}\right\rangle=-{\rm tr}\left\langle\varphi_{1}^{2}\right\rangle$. Since ${\rm tr}\left\langle\varphi_{1}^{2}\right\rangle\neq 0$ (see the beginning of section 3), this tries to make ${\rm tr}\left\langle\varphi_{2}^{2}\right\rangle\neq 0$ for the vacuum solution.
@@
In addition, the $O_{0}$ part of the quantum potential also generates a Majorana mass term along with its scalar mass counterpart:
$\tfrac{4}{9}\bar{O}_{0}\left(\frac{\theta\tilde{\nu}_{W}^{\left(1\right)}}{\bar{\tilde{\phi}}_{21}}+\frac{\theta\nu_{W}^{\left(2\right)}}{\bar{\phi}_{22}}+\frac{\theta\tilde{\nu}_{W}^{\left(2\right)}}{\bar{\tilde{\phi}}_{22}}+\frac{\theta\nu_{W}^{\left(1\right)}}{\bar{\phi}_{21}}\right)^{2}$ (A.11)
$+$ $\tfrac{4}{9}\bar{O}_{0}\theta^{2}\left(\frac{(\tilde{f}_{21})_{1}}{\bar{\tilde{\phi}}_{21}}+\frac{(f_{22})_{2}}{\bar{\phi}_{22}}\right)\left(\frac{(\tilde{\phi}_{21})_{1}}{\bar{\tilde{\phi}}_{21}}+\frac{(\phi_{22})_{2}}{\bar{\phi}_{22}}\right)+...$
Since the vevs $\bar{\tilde{\phi}}_{21}$ and $\bar{\phi}_{22}$ are very small, the first line generates a very large Majorana mass $\tilde{M}_{\nu}$. The second line generates the corresponding scalar mass, where the notation $(\tilde{\phi}_{21})_{1}$ refers only to the first SU(3) component of that scalar.
Due to the relative magnitudes of the vevs, it is a good approximation to treat the above Majorana neutrino as being primarily for the combination $\tilde{\nu}_{M}=\cos\theta_{12}\tilde{\nu}_{W}^{\left(1\right)}+\sin\theta_{12}\nu_{W}^{\left(2\right)}$, where $\tan\theta_{12}=\bar{\tilde{\phi}}_{21}/\bar{\phi}_{22}$.
@@
Below the unification scale, it is possible to consider an effective theory where the $m=2$ SU(3) group has been broken down to SU(2). From section 4 of this paper, below the scale of $\tilde{M}_{\nu}$ (the heavy Majorana neutrino), the SU(2) group has a value of $b_{2}=-\tfrac{1}{6}$ for its running coupling. Following the logic earlier in this appendix, the following “meson” factors are gauge invariant to the SU(2), $U(1)_{Y}$ and $U(1)_{Y^{\prime}}$ groups and also satisfy the axial anomaly selection rule:
$M_{mm^{\prime}FF^{\prime}}^{\prime\left(n\right)}=\Lambda^{-1/18}_{\rm SU(2)}\tilde{Q}_{mF}T_{2}^{0}\Phi^{n}T_{2}^{0}Q_{m^{\prime}F^{\prime}}.$ (A.14)
Due to the factors of $T_{2}^{0}$ (from eq (1.7)), the above terms pick out leptons and scalars that only experience the SU(2) interaction, not the SU(3) strong interaction. The factor of $1/18$ is due to $b_{2}$ in the effective theory being $1/18$ the value of $b_{3}$ of eq (A.2) and the fact that all three flavors of leptons are still in the effective theory.
Since SU(2) is a self-conjugate group, there are additional terms in the SU(2) sector. Consider the following definitions for the $F=1,2$ fields in that sector:
$Q^{\prime}_{F}$ $=\left(\begin{array}[]{c}(\phi_{2F})_{1}+(\theta^{2}f_{2F})_{1}+\sqrt{2}\theta(\psi_{2F})_{1}\\ (\phi_{2F})_{2}+(\theta^{2}f_{2F})_{2}+\sqrt{2}\theta(\psi_{2F})_{2}\\ (\tilde{\phi}_{2F})_{3}+(\theta^{2}\tilde{f}_{2F})_{3}+\sqrt{2}\theta(\tilde{\psi}_{2F})_{3}\end{array}\right)$ (A.15)
$\tilde{Q}^{\prime T}_{F}$ $=\left(\begin{array}[]{c}(\tilde{\phi}_{2F})_{1}+(\theta^{2}\tilde{f}_{2F})_{1}+\sqrt{2}\theta(\tilde{\psi}_{2F})_{1}\\ (\tilde{\phi}_{2F})_{2}+(\theta^{2}\tilde{f}_{2F})_{2}+\sqrt{2}\theta(\tilde{\psi}_{2F})_{2}\\ (\phi_{2F})_{3}+(\theta^{2}f_{2F})_{3}+\sqrt{2}\theta(\psi_{2F})_{3}\end{array}\right),$
where the lower index outside of parentheses in each element refers to the original fundamental SU(3) index of that field. In $Q^{\prime}_{F}$ and $\tilde{Q}^{\prime}_{F}$, the third elements (the SU(2) singlets) have been changed to conjugate fields. The following terms are gauge invariant to the SU(2), $U(1)_{Y}$ and $U(1)_{Y}^{\prime}$ groups while also satisfying the axial anomaly selection rule:
$\Lambda^{-1/18}_{\rm SU(2)}\left(\tilde{Q}_{F}\varphi_{2}T^{\prime 2}\tilde{Q}^{\prime T}_{F}+\tilde{Q}^{\prime}_{F}T^{\prime 2}\varphi_{2}^{T}\tilde{Q}^{T}_{F}\right)$ (A.16)
$\Lambda^{-1/18}_{\rm SU(2)}\left(Q^{\prime T}_{F}T^{\prime 2}\varphi_{2}Q_{F}+Q^{T}_{F}\varphi_{2}^{T}T^{\prime 2}Q^{\prime}_{F}\right),$
where the 3$\times$3 matrix $T^{\prime 2}$ is like $t^{2}$ (half the Pauli matrix), except with $\tfrac{1}{2}$ in the “33” element. Also $Q_{F}$ (with no prime) in the above equation represents the bottom 3 components of the 6-vector $Q_{1F}+Q_{2F}$, and $\tilde{Q}_{F}$ is defined the same way.
Following similar logic to that of eqs (A.10) and (A.11), the superpotential terms of eq (A.16) will lead to flavor-1 charged lepton masses proportional to $c/\bar{\tilde{\phi}}_{21}^{2}$ and $1/\bar{\phi}_{21}^{2}$ for the Omega lepton and electron, respectively. In other words, it is proposed that the vevs generate the ratio
$m_{\Omega}:m_{e}=c/\bar{\tilde{\phi}}_{21}^{2}:1/\bar{\phi}_{21}^{2}\sim 10^{6}-10^{9}:1$ (A.17)
for mass terms $m_{\Omega}\tilde{e}_{W}^{(1)}e^{(1)}$ and $m_{e}\tilde{e}^{(1)}e_{W}^{(1)}$. The factor of $c$ is from eq (3.4). The range of multiples is to accommodate the wide range of possible Omega lepton masses described in section 3.
Due to the structure of eq (A.15), for each pair of two 2-component leptons that combine to form a Dirac mass, there are two scalar fields with that same mass contribution. In particular, the same mechanism that generates a mass of $m_{\Omega}$ for the Omega lepton generates that same mass contribution to the scalar fields $(\phi_{21})_{3}$ and $(\tilde{\phi}_{21})_{2}$. The total mass of these scalar fields can also include contributions from the scalar mass parameters $m_{21}$ and $\tilde{m}_{21}$ in the Lagrangian. Since $m_{21}$ should be significantly smaller than $m_{\Omega}$, the mass of $(\phi_{21})_{3}$ should be similar to that of the Omega lepton.

### v8

@@ 5.7 Scalars
As discussed in section 3, in an $N=2$ Super Yang-Mills (SYM) theory where the adjoint superfield has a tree-level mass, the adjoint scalars become massless color monopoles. The monopoles condense and cause confinement [16, 17, 14, 13]. Since the $m=1$ sector of this theory has a very similar moduli space, it was proposed in section 3 that the octet of $m=1$ SU(3) adjoint scalars of this theory become monopoles, condense, and cause color confinement in the same way. Above the confinement scale, these scalars also participate in the processes mentioned in the Hadronic Cross Section subsection above.
The model’s remaining neutral scalars that have not yet been discussed are $(\phi_{22})_{3}$, $\varphi_{1}^{0}$, $\varphi_{2}^{0}$ and $\varphi_{2}^{8}$. None of these interact with any gauge bosons below the unification scale. The last three of these can be re-expressed as scalars $\varphi_{5}$, $\varphi_{6}$ and $\varphi_{5^{\prime}}$ with the following group structures: $T_{5}=\tfrac{1}{\sqrt{12}}(1,1,1,1,1,-1)$, $T_{6}=\tfrac{1}{\sqrt{60}}(1,1,1,1,1,5)$, and $T_{5^{\prime}}=\sqrt{\tfrac{3}{20}}(\tfrac{2}{3},\tfrac{2}{3},\tfrac{2}{3},-1,-1,0)$.
The scalar $\varphi_{5^{\prime}}$ has supergauge interactions (with gaugino quarks). For this reason, it is assumed that radiative corrections cause it to have a mass closer to the others in the $\varphi_{2}$ octet: $\sim$90-150 GeV. It has supergauge and superpotential decays to $q\bar{q}$.
The scalars $(\phi_{22})_{3}$, $\varphi_{5}$ and $\varphi_{6}$ have no interactions with gaugino quarks, so they have no supergauge interactions at all. For this reason, radiative corrections for these scalars may be smaller, and consequently their masses may be smaller than the others. $\varphi_{6}$ has a stronger interaction with leptons than $\varphi_{5}$, and $(\phi_{22})_{3}$ has no interaction with leptons. The strong superpotential coupling that leads to the heavy Omega lepton masses predicted by this model may cause $\varphi_{6}$ to have a larger mass than $\varphi_{5}$, and $(\phi_{22})_{3}$. It is assumed that $\varphi_{6}$ has a mass of many tens of GeV, but $\varphi_{5}$ has a much smaller mass, and $(\phi_{22})_{3}$ has the smallest mass.
In the Seventh Quark subsection above, it was suggested that $\varphi_{5}$ plays the dominant role in $f$-quark hadron decay. The primary production mechanism for these scalars is via the prompt decay of the heavy Omega leptons to the known three flavors of light leptons.
With their small masses and very few interactions, both $\varphi_{5}$ and $(\phi_{22})_{3}$ could be stable particles and interesting dark matter candidates.
Some of the arguments above rely on quantum corrections that have been qualitatively discussed, but not calculated in detail. Future work would be needed to ensure that the quantum corrections of this model are consistent with the picture painted above.
@@ Discussion
In sections 1, 2, 4 and the appendix, a model is presented that has a number of attractive theoretical features. For example, it is holomorphic, invariant to local superspace gauge transformations, supports coupling constant unification, and is similar at the unification scale to a theory that has been shown to be free of quadratic divergences to at least two loops [3].
@@
But even without that work, the model provides a couple of interesting explanations and makes a number of predictions. The model provides explanations for the mechanisms of confinement and neutrino oscillations. The model predicts three additional charged leptons, a seventh quark (without an eighth), a $Z^{\prime}$ boson, right-handed quark interactions with the $W$ boson, and dozens of additional scalar particles. So far, these predictions do not appear to be ruled out by existing data; in fact, they could provide new physics explanations for many of the anomalies discussed in [4].
@@ Appendix A Quantum Effective potential
The U(3)xU(3) theory of this paper has Abelian groups, so it not entirely asymptotically free. However, in this appendix it is implicitly assumed that this theory is an effective theory of a more general asymptotically free theory, applicable above the unification scale. With that assumption in hand, this appendix makes arguments similar to those used for SQCD and proposes general features of nonperturbative superpotential terms.
@@
Since they are in the $(3,\tilde{3})$ and $(\tilde{3},3)$ representations, the gauginos and the fermions from the twisted adjoint superfield contribute zero to both $T^{0}_{V_{R}+}$ and $T^{0}_{A_{R}+}$. In addition, for the second term of (A.4) they generate a factor of $f^{bad}d^{cad}$ which vanishes. Therefore, those fermions do not contribute to the axial anomaly.
@@
The anomaly-generating axial U(1) charges of the superfields $Q_{mF}$, $\tilde{Q}_{mF}$, and $\Phi$ are 1, 1, and 0, respectively. Therefore, the following factors are invariant to the U(3)$\times$×U(3) local gauge symmetry and obey the axial selection rule:
@@
where $n$ is a non-negative integer. In the second “meson” configuration above, the $m$ and $F$ indices are restricted to combinations that have the same charge for $V_{+}^{0}$ in eq (1.17). For example, $M_{2213}^{\left(n\right)}=0$ (since $q_{21}\neq q_{23}$), but $M_{2113}^{\left(n\right)}\neq 0$ (since $q_{21}=q_{13}$).
@@
In SQCD, the allowed nonperturbatively generated terms also have to respect flavor symmetries. In this model, that is complicated for quantum-generated mass terms since $Q_{1F}$ and $\tilde{Q}_{1F}$ have a $3\times 3$ flavor symmetry (since $q_{11}=q_{12}=q_{13}$), but $Q_{2F}$ and $\tilde{Q}_{2F}$ only have a $1\oplus 2\times 2$ symmetry (since $q_{21}\neq q_{22}=q_{23}$). The flavor symmetry will be incorporated as follows: First, one of the $m=1$ flavors will be identified with the $m=2$ flavor $F=1$ and terms involving flavor determinants will be constructed. Then the other two $m=1$ flavors will be identified with the $m=2$ flavor $F=1$ and analogous terms will be constructed.
Identifying the $m=1$ flavor $F=1$ with the $m=2$ flavor $F=1$, the following combination chiral fields and meson operators can be constructed:
$Q_{F}=Q_{1F}+Q_{2F}\hskip 28.45274pt\tilde{Q}_{F}=\tilde{Q}_{1F}+\tilde{Q}_{2F}$ (A.7)
$M_{FF^{\prime}}^{(n)}=\Lambda^{-1}\tilde{Q}_{F}\Phi^{n}Q_{F^{\prime}}\,.$
In the meson factors, any terms that are not gauge invariant are set to zero. For example,
$M_{12}^{(0)}=\Lambda^{-1}(\tilde{Q}_{11}Q_{12}+\tilde{Q}_{21}Q_{12})\,.$ (A.8)
The terms $\tilde{Q}_{11}Q_{22}=\tilde{Q}_{21}Q_{22}=0$ inside of $M_{12}^{(0)}$ since $q_{21}=q_{11}\neq q_{22}$.
The following determinant is then defined
$\Delta^{(pqr)}={\rm det}\left|\begin{array}[]{ccc}{M_{11}^{(p)}}&{M_{12}^{(q)}}&{M_{13}^{(r)}}\\ {M_{21}^{(p)}}&{M_{22}^{(q)}}&{M_{23}^{(r)}}\\ {M_{31}^{(p)}}&{M_{32}^{(q)}}&{M_{33}^{(r)}}\end{array}\right|\,$ (A.9)
Following symmetry arguments similar to those used for SQCD, quantum interactions will generate low energy effective superpotential terms with mass dimension 3 and positive powers of $\Lambda$ that are constructed from the above determinants as well as factors of ${\rm Tr}\left(\Phi^{r}\right)$, and other gauge invariant factors such as those mentioned later in this Appendix.
@@
$O_{0}=\left({\rm Tr}\left(\Phi^{2}\right)\right)^{2}\left(\Delta^{(000)}\right)^{-1/3}\,.$ (A.10)
The scalar potential can be derived from the superpotential terms like the one above by taking derivatives to extract “f terms” proportional to $\theta^{2}f$ and inserting them into the f-term part of eq (2.1). The quantum vacuum can be found by minimizing the vev of the scalar potential after including these quantum contributions.
In this process, vevs of scalar potential terms derived from $O_{0}$ will be proportional to
@@
In section 2, it was noted that if the model includes large tree-level masses $\tilde{m}_{21}$ and $\tilde{m}_{22}$, then the vevs $\bar{\tilde{\phi}}_{21}$ and $\bar{\tilde{\phi}}_{22}$ vanish clasically. However, in order to stabilize expressions such as the one above, it is assumed that $\bar{\tilde{\phi}}_{21}$ and $\bar{\tilde{\phi}}_{22}$ acquire small vevs quantum mechanically. This is mentioned in section 3. As shown below, those quantum-generated small vevs lead to a very large right-chiral neutrino Majorana mass term that allows the model to reproduce observed neutrino masses and mixing.
The classical scalar potential only restricts the difference $\bar{\phi}_{23}^{2}-\bar{\tilde{\phi}}_{23}^{2}$. However, it can be seen from eq (A.11) that the quantum scalar potential will try to make each of these vevs go to infinity while maintaining the difference. A mechanism like that is what causes Supersymmetric QCD with fewer flavors than colors to not have a vacuum solution. But as described in section 2, a nonzero tree-level or quantum-generated superpotential coupling provides a small counterbalancing effect that stops $\bar{\phi}_{23}$ from becoming infinite.
Minimizing the quantum scalar potential generates the adjoint vevs in $\left\langle\varphi_{2}\right\rangle$ of eq (3.4) to reduce ${\rm tr}\left\langle\varphi_{2}^{2}\right\rangle=-{\rm tr}\left\langle\varphi_{1}^{2}\right\rangle$ in scalar potential contributions proportional to $\bar{O}_{0}$. Since ${\rm tr}\left\langle\varphi_{1}^{2}\right\rangle\neq 0$ (see the beginning of section 3), minimization makes ${\rm tr}\left\langle\varphi_{2}^{2}\right\rangle\neq 0$ for the vacuum solution.
@@
The quantum potential also generates a Majorana mass term along with its scalar mass counterpart:
$\frac{4\theta^{2}\bar{O}_{0}}{9\bar{\tilde{\phi}}_{21}^{2}}\left(2(\tilde{f}_{21})_{1}(\tilde{\phi}_{21})_{1}-\tilde{\nu}_{W}^{(1)}\tilde{\nu}_{W}^{(1)}\right)\,,$ (A.13)
where terms involving $1/\bar{\phi}_{21}$ have been ignored. Since the vev $\bar{\tilde{\phi}}_{21}$ is very small, the second term generates a very large neutrino Majorana mass. The first term generates the corresponding scalar mass, where the notation $(\tilde{\phi}_{21})_{1}$ refers only to the first SU(3) component of that scalar.
The quantum-generated mass terms in eqs (A.12) and (A.13) were constructed in the context of the $m=1$ flavor $F=1$ being identified with the $m=2$ flavor $F=1$. Due to the $3\times 3$ flavor symmetry in the $m=1$ sector, there are also analogous terms where the $m=1$ flavors $F=2$ or $F=3$ are identified with the $m=2$ flavor $F=1$. With those included, the neutrino Majorana mass term becomes
$\frac{4\theta^{2}\bar{O}_{0}}{9\bar{\tilde{\phi}}_{21}^{2}}\sum_{F}\tilde{\nu}_{W}^{(F)}\tilde{\nu}_{W}^{(F)}\,.$ (A.14)
In other words, all three right-handed neutrinos have the same very large Majorana mass, $\tilde{M}_{\nu}$. For $m=2$ scalars, there is no analogous flavor rotation since the $m=2$ sector only has $1\oplus 2\times 2$ flavor symmetry. So only $(\tilde{\phi}_{21})_{1}$ gets a very large mass by the above mechanism.
Under the above $m=1$ flavor rotation, the $m=1$ mass terms of eq (A.12) become
$-\tfrac{1}{3}\bar{O}_{0}\left(\frac{1}{\bar{\tilde{\phi}}_{21}\bar{\phi}_{21}}+\frac{1}{\bar{\tilde{\phi}}_{22}\bar{\phi}_{22}}\right)\sum_{F}\tilde{Q}_{1F}Q_{1F}\,.$ (A.15)
The same mechanism that generates large Majorana neutrino masses can also generate an effective coupling important for scalar-mediated quark decays. Just as quantum effects generate $O_{0}$ in eq (A.10), they will also generate an $O_{2}$ that replaces $\Delta^{(000)}$ in $O_{0}$ with $\Delta^{(200)}+\Delta^{(020)}+\Delta^{(002)}$, while changing $({\rm Tr}(\Phi^{2}))^{2}\to({\rm Tr}(\Phi^{2}))^{7/3}$. This generates the following quantum superpotential term:
$-i\frac{\sqrt{2}\theta^{2}\bar{O}_{c}}{3c\bar{\varphi}_{2}^{2}\bar{\tilde{\phi}}_{21}}\left(\tilde{u}^{(1)}\varphi_{1}u_{W}^{A}\right)\,.$ (A.16)
Just as for the Majorana neutrino masses, since $\bar{\tilde{\phi}}_{21}$ is very small, this quantum-generated coupling is large. This coupling is used in section 5 when showing how the model can reproduce $Z$ pole observables.